# Homogenization with Guaranteed Bounds via Primal-Dual Physically Informed Neural Networks

### v1

@@ 4.2.2 VPINN with neural network test functions (VNPINN)
From the presented results, we conclude that the large networks (1,801 and 15,601 parameters) gain around 1% precision of the primal estimate for each doubling of the test basis, and up to 4% in the dual estimate. The point of diminishing returns is reached around 400 test functions, where the only improvement is in the primal solution by the largest network. This is explained by the overall slowdown in training for such a large number of test functions (Fig. 17), compared to the smaller test basis (Fig. 16).
@@ 5 Discussion
The relative error of the final primal and dual estimates for all networks is summarised in Tab. 2. Tab. 3 then lists the computed guaranteed error bounds from the piecewise linear approximations of the solutions.
@@
As shown in Table 2, the strong-form PINNs demonstrate superior convergence for smoothly approximated materials with $\varepsilon\in\{1/10,1/20,1/30\}$, achieving relative errors below 0.5% for the largest networks in the primal formulation. However, this performance degrades catastrophically when $\varepsilon=1/40$, with relative errors exceeding 10% even for the 15,601-parameter network, indicating a critical failure mode when the material transition becomes too sharp.
Among the variational approaches, VNPINNs with neural network test functions outperform VSPINNs for most configurations examined. The VSPINNs exhibit non-monotonic behavior with respect to network size, as evidenced by the degraded performance of the 15,601-parameter network with 126 test functions (7.054% error) compared to smaller networks. This suggests overfitting to integration errors when high-frequency spectral basis functions are employed with insufficient quadrature precision. In contrast, VNPINNs show more consistent improvement with increasing test functions, achieving their best performance with 400 test functions for the largest networks, reaching 0.460% relative error in the primal formulation.
The guaranteed bounds presented in Tab. 3 are close to the estimates, with the projection onto piecewise linear finite elements introducing minimal additional error (typically less than 0.5%). Notably, the bounds maintain the theoretical ordering property of Eq. (9) for all solutions, providing a reliable indicator of solution quality.
@@ 6 Conclusion
We examined the effectiveness of dual formulations within two primary strategies for addressing PINN limitations with discontinuous materials: smooth approximation of material coefficients in strong-form PINNs, and variational formulations with both spectral and neural network test function bases. Our findings reveal fundamental trade-offs between these approaches. Strong-form PINNs achieve superior accuracy for sufficiently smooth material approximations but require careful calibration of the smoothing parameter to avoid catastrophic failure. VPINNs, though computationally more demanding and sensitive to test function selection, offer greater robustness by directly accommodating discontinuous material properties.
@@
LG: Conceptualization, Methodology, Software, Investigation, Writing – Original Draft, Visualization, Funding Acquisition; IP: Methodology, Writing – Review & Editing, Supervision, Funding Acquisition; MD: Methodology, Writing – Review & Editing, Supervision; OR: Methodology, Writing – Review & Editing, Supervision.
@@ Acknowledgments
LG, IP, and MD acknowledge funding by the European Union under the project ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590). LG also acknowledges the Student Grant Competition of the Czech Technical University in Prague (project no. SGS25/002/ OHK1/1T/11).

### v2

@@ 4.1.2 Weak form
From the presented results, we conclude that the large networks (1 8011\,801 and 15 60115\,601 parameters) gain around 1% precision of the primal estimate for each doubling of the test basis, and up to 4% in the dual estimate. The point of diminishing returns is reached around 400 test functions, where the only improvement is in the primal solution by the largest network. This is explained by the overall slowdown in training for such a large number of test functions (Fig. 17), compared to the smaller test basis (Fig. 16).
4.2 Multiple Inclusions
The performance of strong-form PINNs in the primal formulation was thoroughly tested by JIANG2023115972 on multi-inclusion composites with varying volume fraction, where it was shown to achieve the same precision in the homogenized parameters as FEM. However, the aforementioned study did not consider the error introduced by the material smoothing. Here, we assume a composite with four randomly distributed equi-sized inclusions in a $2\pi\times 2\pi$ periodic cell, where $\gamma_{\mathrm{mat}}=0.1$ and $\gamma_{\mathrm{inc}}=1$ are thermal conductivities of the matrix and the inclusions, respectively, see Fig. 20. For the PINN solution, we assume the smooth approximation of the conductivity function as
[…]
where $\bm{x}^{c}_{i}$ are the centers of the inclusions, $r$ is the radius of the inclusions, and $\varepsilon>0$ controls the sharpness of the transition. Since the previous example with a single inclusion showed that there is an optimal $\varepsilon$ that minimizes the material approximation error while still allowing the PINN to converge to a high-quality solution, we only test the PINNs on one such $\varepsilon$, which we found to be roughly equal to 0.06 of the inclusion radius $0.2\pi$, or $0.012\pi$.
Figure 20: Smooth approximation (a) $\bm{A}_{s,\varepsilon}(\bm{x})$ of the material with multiple inclusions (b)
As there is no known analytical solution for this problem, we compare the neural network solutions only to the FEM benchmark computed on a $200\times 200$ regular mesh, which provides the upper and lower bounds within 0.35% of each other, see Fig. 21. We then test two larger configurations of the neural networks (Tab. 2) in the primal and dual formulation as PINNs with the smooth material approximations and VNPINNs with 400 test functions. The collocation points are selected on the $200\times 200$ regular grid, which is then used for trapezoidal integration. All networks were trained for 100 000100\,000 epochs using the Adam optimizer and the initial learning rate of $10^{-5}$.
Figure 21: Benchmark primal and dual solutions for $\bm{\xi}=[1,0]^{\rm{T}}$, obtained with linear triangular finite elements, using discretization into $200\times 200$ unique DoFs.
№ neurons, periodic layer № neurons, per residual layer № residual layers № parameters
50 50 5 15 60115\,601
70 70 6 35 21135\,211
Table 2: Configurations of the considered neural networks for both primal versus dual and strong versus weak formulations, for the multiple inclusions problem.
The estimates and bounds of the homogenized parameter $A_{1,1}^{*}$ from the primal and dual solutions for PINNs and VPINNs are shown in Fig. 22, normalized by the dual FEM bound. Both 15 60115\,601 and 35 21135\,211-parameter PINN converged to virtually the same parameter estimate in the primal and dual formulation. However, when the material approximation error is considered, PINNs provide an estimate around 2% higher than the FEM benchmark, while the calculated bounds on the exact material distribution are up to 1.2% away from the FEM benchmark. Even with the material approximation error, PINNs have achieved a closer result to the benchmark than the VNPINNs that have been trained on the exact material distribution, whose primal bounds and estimates come to around 2.5% higher than the benchmark, and dual bounds and estimates are 0.5% lower. The PINN and VNPINN solutions and their residuals can be compared in Figs. 23 and  24. Interestingly, unlike in the previous example, the dual formulation provided a more accurate estimate. One possible explanation could lie in the opposite contrast ratios of selected conductivities in the matrix and inclusion in the two examples. The dependence on the relative conductivity might be corroborated by the observation that the strong-form residuals of the dual VPINNs are significantly higher for the less conductive (i.e., more resistant) square inclusion in comparison to the matrix in Fig. 19, while more conductive circular inclusions have higher residuals in the primal formulation in Fig. 24. Strong-form PINNs typically only have significantly higher residuals at the interface of the materials. The behavior of VPINNs for different material contrasts should be explored in a future study.
Figure 22: Precision of the primal and dual estimates of the PINN and VNPINN depending on the number of network parameters.
Figure 23: Primal solutions (top) and the residuals (bottom) of PINNs and VPINNs for the multiple inclusion problem.
Figure 24: Dual solutions (top) and the residuals (bottom) of PINNs and VPINNs for the multiple inclusion problem.
The training of PINNs and VPINNs is shown in Fig. 25 and Fig. 26. The VPINN estimates stopped improving after about 20 00020\,000 epochs, which suggests the overfitting to the test functions, while PINN estimates achieved a similar precision slightly slower (around 30 00030\,000 epochs) but continued to improve significantly. However, considering that material smoothing puts a hard limit on the guaranteed bounds, the utility of such a prolonged training of PINNs should be assessed.
Figure 25: Training loss of the primal and dual solutions, and the gap between the estimates, PINN, multiple inclusion problem.
Figure 26: Training loss of the primal and dual solutions, and the gap between the estimates, VNPINN, multiple inclusion problem.
@@ 5 Discussion
The relative error of the final primal and dual estimates for all networks of the single inclusion experiment (Section 4.1) is summarised in Tab. 3. Tab. 4 then lists the computed guaranteed error bounds from the piecewise linear approximations of the solutions.
@@
As shown in Tab. 4, the strong-form PINNs demonstrate superior convergence for smoothly approximated materials with $\varepsilon\in\{1/10,1/20,1/30\}$, achieving relative errors below 0.5% for the largest networks in the primal formulation. However, this performance degrades catastrophically when $\varepsilon=1/40$, with relative errors exceeding 10% even for the 15 60115\,601-parameter network, indicating a critical failure mode when the material transition becomes too sharp.
Among the variational approaches, VNPINNs with neural network test functions outperform VSPINNs for most configurations examined. The VSPINNs exhibit non-monotonic behavior with respect to network size, as evidenced by the degraded performance of the 15,601-parameter network with 126 test functions (7.054% error) compared to smaller networks. This suggests overfitting to integration errors when high-frequency spectral basis functions are employed with insufficient quadrature precision. In contrast, VNPINNs show more consistent improvement with increasing the number of test functions, achieving their best performance with 400 test functions for the largest networks, reaching 0.460% relative error in the primal formulation.
The guaranteed bounds presented in Tab. 4 are close to the estimates, with the projection onto piecewise linear finite elements introducing minimal additional error (typically less than 0.5%). With precise integration, the bounds calculated from the projected solution maintain the theoretical ordering property of Eq. (18) for all solutions, providing a reliable indicator of solution quality. Unfortunately, this requires a careful meshing of the geometry, which is not always feasible and diminishes one of the primary benefits of PINNs. However, in practice, a reasonably precise numerical integration typically produces reliable results as well, as demonstrated by the primary estimates that can be computed fast during the training of the network.
For the multiple inclusions example, we similarly conclude that strong-form PINNs can perform extremely well for smooth geometries, but cannot match even simple piece-wise linear FEM for discontinuous materials. The quality of the test basis remains the main limiting factor for VPINNs, and the number of required test functions depends on the complexity of the geometry as well.
@@ 6 Conclusion
We examined the effectiveness of dual formulations within two primary strategies for addressing PINN limitations with discontinuous materials: (i) smooth approximation of material coefficients in strong-form PINNs, and (ii) variational formulations with both spectral and neural network test function bases. Our findings reveal fundamental trade-offs between these approaches. Strong-form PINNs achieve superior accuracy for sufficiently smooth material approximations but require careful calibration of the smoothing parameter to avoid catastrophic(and unidentifiable without the duality gap calculation ) failure. VPINNs, though computationally more demanding and sensitive to test function selection, offer greater robustness by directly accommodating discontinuous material properties.
@@
LG: Conceptualization, Methodology, Software, Investigation, Writing – Original Draft, Visualization, Funding Acquisition; OR: Methodology, Writing – Review & Editing, Supervision; MD: Methodology, Writing – Review & Editing, Supervision; IP: Methodology, Writing – Review & Editing, Supervision, Funding Acquisition.
@@ Acknowledgments
LG and IP acknowledge funding by the European Union under the project ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590). MD’s work was supported by the Czech Science Foundation, project No. 24-11845S. LG also acknowledges the Student Grant Competition of the Czech Technical University in Prague (project no. SGS25/002/ OHK1/1T/11).